OpenStudy (anonymous):

1. A student collected the following data for a fixed volume of gas: Temperature (⁰C) Pressure (mm of Hg) 10 726 20 750 40 800 70 880 100 960 150 ??? Fill in the missing data point. Show all calculations leading to an answer.

OpenStudy (jtfan2000):

100 / 960 = 150 / x 960 x 150 = 100x ??? = 1440 mm of Hg